OpenStudy (anonymous):

Marvin was completing the square, and his work is shown below. Identify the line where he made his mistake. f(x) = 2x^2 − 8x + 5 Line 1: f(x) = 2(x^2 − 8x) + 5 Line 2: f(x) = 2(x^2 − 8x + 16) + 5 − 32 Line 3: f(x) = 2(x − 4)^2 − 27 Line 1 Line 2 Line 3 He did not make any mistakes I think line 2, amirite?

OpenStudy (anonymous):

yes, he didn't factor the 2 out correctly

OpenStudy (dinamix):

f(x) =2x^2-16x+5 this equation i think the answer is line 3

OpenStudy (anonymous):

oh wait, sorry. That factoring mistake is on line 1, not 2.

OpenStudy (dinamix):

yup this @peachpi and he didnt right the correct must will f(x) = 2x^2-16x+5 and the aswer line 3

OpenStudy (dinamix):

2(x^2-4) @peachpi

OpenStudy (anonymous):

Y'all lost me

OpenStudy (anonymous):

\(2x^2-8x\) = \(2(x^2-4x)\) is what I meant

OpenStudy (anonymous):

they're asking you to find the mistake. The mistake is that the factoring on line 1 was wrong.

OpenStudy (dinamix):

all lines is wrong @peachpi

OpenStudy (anonymous):

yes, they're all wrong. But the mistake first happened on line 1

OpenStudy (anonymous):

Oh, alright, I see where you guys are coming from

OpenStudy (dinamix):

i think he made mistake in f(x) = 2x^2-8x+5 and should be 2x^2-16x+5 peachpi

OpenStudy (dinamix):

and the lines will be all correct

OpenStudy (dinamix):

did understand me @peachpi

OpenStudy (anonymous):

No. I don't know where you're getting that. The two expressions you typed aren't equal. This problem is about completing the square, which is basically rewriting the same thing in another way. You can't just change -8x to -16x because it changes the value of the expression.

OpenStudy (dinamix):

yup he did write it correct thats problem

OpenStudy (dinamix):

u didnt understand me if the function f(x) = 2x^2-16x+5 we all lines correct right ?

OpenStudy (dinamix):

find*

OpenStudy (anonymous):

No, because changing -8 to -16 changes the function. This is the correct solution to the problem, with no mistakes \(f(x)=2x^2-8x+5\) \(f(x)=2(x^2-4x)+5\) \(f(x)=2(x^2-4x+4)+5-8\) \(f(x)=2(x-2)^2-3\)

OpenStudy (anonymous):

Is he getting confused with the actual function? Hes really confusing me

OpenStudy (anonymous):

yeah I think so. If you compare what I have with the solution in the problem statement, you can see the mistake on the first line

OpenStudy (anonymous):

He didn't divide the 8x by the 2 you're factoring?

OpenStudy (dinamix):

@gigirained are really this function f(x) = 2x^2-8x+5 are u sure its -8x not -16x ?

OpenStudy (anonymous):

yes correct @gigirained

OpenStudy (anonymous):

Positive its -8x

OpenStudy (anonymous):

@dinamix I see what you're doing. The objective of the problem is to find the mistake, if there is one. You're finding a function that matches the steps provided. So yeah if the initial function had -16x, then the work would be correct.

OpenStudy (dinamix):

ok he did mistake , didnt divide 8x by 2

OpenStudy (dinamix):

yup @peachpi this why i check if write function correct or no

OpenStudy (anonymous):

Thanks you guys, @peachpi and @dinamix